Tanya is selling her old car for $6,933. If a prospective buyer offers to make yearly payments for the next 5 years, what should Tanya set these payments to if the correct discount rate is 0.11?
Answer #1

Value of car =  P = $6933

Interest Rate = r = 0.11

Number of payment periods = n = 5

Let annual payments made be X

Hence, the sum of present value of annual payments must be equal to the value of the loan amount

=> X/(1+r) + X/(1+r)2 +....+ X/(1+r)N = P

=> X[1- (1+r)-N]/r = P

=> X = rP(1+r)N/[(1+r)N-1]

Hence, MoAnnualnthly Payments =  rP(1+r)N/[(1+r)N-1]

= 6933*( 0.11)*(1+ 0.11)5/((1+ 0.11)5-1) = $1875.86